OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. (Do not type the variable in the blank.) 4x2 + ___x + 49

OpenStudy (anonymous):

so its 14

OpenStudy (asnaseer):

to be a perfect square, it has to be of the form \((ax+b)^2\). So if we expand this we get:\[(ax+b)^2=a^2x^2+2abx+b^2\]Compare this to your trinomial:\[4x^2+?x+49\]to determine the values for a and b.

OpenStudy (cwrw238):

sorry - i didn't notice the 4 before the x2 - i need new glasses asnaseer as given you the method hint : a^2 = 4 ( comparing coefficients of x^2)